OpenStudy (anonymous):

medals!!!! solve 5+√x+2=8+√x-7

OpenStudy (anonymous):

@ganeshie8

OpenStudy (jhannybean):

\[5+\sqrt{x+2} =8+\sqrt{x-7}\]Square both sides \[(5+\sqrt{x+2})^2=(8+\sqrt{x-7})^2\] what do you get ?

OpenStudy (anonymous):

13√x+3^2 so 13 √9

OpenStudy (jhannybean):

Hmm, let me check. \[(5+\sqrt{x+2})^2=(8+\sqrt{x-7})^2\]\[\implies 25 +10\sqrt{x+2}+x+2=64 +16\sqrt{x-7} +x-7\]This is what I got thus far.

OpenStudy (jhannybean):

Simplifies to \[27 +10\sqrt{x+2} = 57+16\sqrt{x-7}\]

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (anonymous):

27+10√x+2?

OpenStudy (jhannybean):

Wait, is it \(\sqrt{x}\) or \(\sqrt{x+2}\)?

OpenStudy (jhannybean):

If it's he first then we're doing it wrong.

OpenStudy (anonymous):

its the second

OpenStudy (jhannybean):

Okay. I think there may be an easier method in solving this. One moment.

OpenStudy (anonymous):

I think you'd want to get rid of the \(5\) or the \(8\) first.

OpenStudy (jhannybean):

You have \[5+\sqrt{x+2}=8+\sqrt{x-7}\]You can isolate one of the radicals.\[\sqrt{x+2} = 3+\sqrt{x-7}\]NOW you square both sides. what we did at first made it a little bit more confusing. \[x+2 = (3+\sqrt{x-7})^2\]\[x+2 = 9+6\sqrt{x-7}+x-7\]Move everything over to the left hand side

OpenStudy (jhannybean):

Move everything over and simplify. \[-6\sqrt{x-7}=0\]divide both sides by -6\[\sqrt{x-7}=0\]square both sides \[x-7=0\]Add -7 to both sides and you get?

OpenStudy (jhannybean):

Let me know if there is a sep you do not understand.

OpenStudy (jhannybean):

step*

OpenStudy (anonymous):

yea that will be great

OpenStudy (jhannybean):

Anything confusing?

OpenStudy (jhannybean):

I practically gave you the answer, all you have to do is solve for x by adding +7 to both sides and you're asking me for the answer?

OpenStudy (anonymous):

x+9?

OpenStudy (anonymous):

then 9√6

OpenStudy (jhannybean):

Go back up to my post where I break down the problem step by step and you will find yoru x value.

OpenStudy (jhannybean):

I do not know where you are getting these numbers.